A NaOH solution is standardized by titration with 0.835 g KHP (monoprotic standard acid, MW = 204.2 g/mol). 23.47 mL of the unknown NaOH solution is needed to titrate the 0.835 g KHP to the endpoint. Calculate the concentration of the NaOH solution.
1
Expert's answer
2015-11-16T09:50:02-0500
The amount of sodium hydroxide is the same as of monoprotic standard acid KHP. So we firstly calculate the amount of it and then from the formula, we can calculate the concentration:
NaOH + KHP = K(Na)P + H2O
n(NaOH)=n(KHP)=m(KHP)/M(KHP)=0.835/204.2=0.0041 mol
V(NaOH)=23.47 mL=0.02347 L, so the concentration is:
c(NaOH)=n(NaOH)/V(NaOH)=0.0041/0.02347=0.174 mol/L

Answer:
c(NaOH)=0.174 mol/L